# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ 4.2.3 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]$
Given $T\in(0,1)$, we have
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]\leq\sigma\sqrt{d^{*}(1-T)}.$
4.2.4 Oracle Inequality
By amalgamating the analyses articulated in Theorem 4.7, Theorem 4.8, and Lemma 4.9, we derive upper bounds for $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},{\widehat{p}}^{*}_{data})]$ through the meticulous determination of a suitable score neural network and the optimal stopping time $T$. The end-to-end convergence rates are shown in the following two theorems.
@@ 4.2.4 Oracle Inequality
Suppose that Assumption 4.3 holds, and the score estimator $\hat{s}$ defined in (6) is structured as introduced in Theorem 4.6. By choosing the maximal step size $\mathop{\text{max}}_{k=0,1,\cdots,N-1}|t_{k+1}-t_{k}|=\mathcal{O}\left(n^{-\frac{2}{d^{*}+3}}+n^{\frac{d^{*}+1}{2(d^{*}+3)}}m^{-\frac{1}{2}}\right)$ and the early stopping time $T=1-(\log{n})^{-\frac{1}{4}}$, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},{\widehat{p}}^{*}_{data})]=\widetilde{\mathcal{O}}\left((\log{n})^{-\frac{1}{8}}+e^{\frac{10{d^{*}}^{5/2}}{\sigma^{2}}\sqrt{\log{n}}}\left(n^{-\frac{1}{d^{*}+3}}+n^{\frac{d^{*}+1}{4(d^{*}+3)}}m^{-\frac{1}{4}}\right)\right).$
In Theorem 4.10, the consistency of the generated samples is established. Specifically, it is demonstrated that $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},{\widehat{p}}^{*}_{data})]\rightarrow{0}$ as $e^{\frac{10{d^{*}}^{5/2}}{\sigma^{2}}\sqrt{\log{n}}}\left(n^{-\frac{1}{d^{*}+3}}+n^{\frac{d^{*}+1}{4(d^{*}+3)}}m^{-\frac{1}{4}}\right)\rightarrow{0}$ and $n\rightarrow{\infty}$. Moreover, this logarithmic rate can be improved with the introduction of the smoothness condition (Assumption 4.4).
Theorem 4.11.
Suppose that Assumption 4.3 and Assumption 4.4 hold, and the score estimator $\hat{s}$ defined in (6) is structured as introduced in Theorem 4.6. By choosing the maximal step size $\mathop{\text{max}}_{k=0,1,\cdots,N-1}|t_{k+1}-t_{k}|=\mathcal{O}\left(n^{-\frac{7}{3(d^{*}+3)}}\right)$ and the early stopping time $T=1-n^{-\frac{1}{3(d^{*}+3)}}$, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},{\widehat{p}}^{*}_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+n^{\frac{3d^{*}+13}{12(d^{*}+3)}}m^{-\frac{1}{4}}\right).$
Moreover, if $m>n^{\frac{d^{*}+5}{d^{*}+3}}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}},{\widehat{p}}^{*}_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}\right).$
@@
Theorem 4.11 rigorously establishes the end-to-end theoretical guarantees for the diffusion model within latent space. This is achieved through the seamless integration of score estimation, EM discretization, and early stop techniques. Furthermore, this analysis is independent of the pre-training theory, rendering our theoretical framework applicable to other diffusion models, irrespective of considerations regarding pre-training encoder-decoder structures. This characteristic underscores the independent validity of our theoretical framework. Importantly, the convergence rate $\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}\right)$ in Theorem 4.11 mitigates the curse of dimensionality of raw data, primarily attributed to the implementation of the encoder component.
@@ 4.3 Main Result
Suppose that Assumptions 4.3-4.5 hold, and the score estimator $\hat{s}$ defined in (6) is structured as introduced in Theorem 4.6. By choosing the maximal step size $\mathop{\text{max}}_{k=0,1,\cdots,N-1}|t_{k+1}-t_{k}|=\mathcal{O}\left(n^{-\frac{7}{3(d^{*}+3)}}\right)$ and the early stopping time $T=1-n^{-\frac{1}{3(d^{*}+3)}}$, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T},p_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+n^{\frac{3d^{*}+13}{12(d^{*}+3)}}m^{-\frac{1}{4}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\mathcal{M}^{-\frac{1}{2(d+2)}}+\delta_{0}^{1/2}\right),$
where $\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T}$ denotes the distribution of $\widehat{\boldsymbol{D}}(\widetilde{\mathbf{x}}_{N})$ in Algorithm 1. Moreover, if $m>n^{\frac{d^{*}+5}{d^{*}+3}}$ and $\mathcal{M}>n^{\frac{d+2}{3(d^{*}+3)}}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T},p_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\delta_{0}^{1/2}\right).$
@@
In Theorem 4.12, the convergence rates of $\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T}$ are established by simultaneously considering pre-training encoder-decoder structures and latent diffusion models. This represents an end-to-end result. Recall that $m$ denotes the sample size of Gaussian noises in the latent space, and $\mathcal{M}$ represents the sample size of pre-training. These two terms can be chosen sufficiently large, as Gaussian noises are readily obtainable, and the pre-training phase can be manipulated separately by resorting to pre-existing large-scale models. Therefore, we can set $m>n^{\frac{d^{*}+5}{d^{*}+3}}$ and $\mathcal{M}>n^{\frac{d+2}{3(d^{*}+3)}}$ to obtain the convergence rate of $\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\delta_{0}^{1/2}\right)$. Additionally, if the distribution shift $\epsilon_{p_{data},\widetilde{p}_{data}}$ and pre-training quantity $\delta_{0}^{1/2}$ are both smaller than $n^{-\frac{1}{6(d^{*}+3)}}$, then the convergence rate reduces to $\widetilde{\mathcal{O}}\left(n^{-\frac{1}{6(d^{*}+3)}}\right)$. This finding circumvents the curse of dimensionality inherent in raw data, thereby demonstrating the effectiveness of latent diffusion models from a theoretical perspective.
@@ 5 Conclusion
In this work, we introduce a novel latent diffusion model rooted in the Schrödinger bridge, facilitated by pre-training an encoder-decoder structure. An SDE, defined over the time interval $[0,1]$, is formulated to effectuate the transformation of the convolution distribution into the encoder target distribution within the latent space. The synergy of pre-training, score estimation, EM discretization, and early stopping enables the derivation of desired samples, manifesting an approximate distribution in accordance with the target distribution. Theoretical analysis encompasses pre-training, score estimation theory, and comprehensive end-to-end convergence rates for the generated samples, quantified in terms of the second-order Wasserstein distance. Our theoretical findings provide solid theoretical guarantees for mainstream diffusion models. The obtained convergence rates predominantly hinge upon the dimension $d^{*}$ of the latent space, which mitigates the curse of dimensionality inherent in raw data. However, the dimension $d^{*}$ is prespecified in practical applications. Therefore, the development of an adaptive approach for selecting it can be regarded as future work. Additionally, extending our theoretical framework to analyze diffusion models based on SDEs, such as OU processes and Langevin SDEs, defined over an infinite time horizon $[0,\infty)$ can also be considered as future work.
@@
In this appendix, we provide the detailed proofs of all lemmas and theorems in this paper. Specifically, in Section A, we establish the Lipschitz continuity of the true score function. Moving on to Section B, we derive the error bound for the pre-training procedure. Subsequently, in Sections C-F, we obtain error bounds for the latent diffusion modeling. Finally, in Section G, we present the proof of our main result.
Appendix A Properties of true score function
In this section, we establish the Lipschitz continuity of the true score function by leveraging the properties inherent in the target distribution. We first derive the Lipschitz constant for $\nabla\log q(\sqrt{1-t}\sigma,\mathbf{x})$ with respect to $\mathbf{x}$ in Lemmas 4.2-4.3. Then we derive the Lipschitz continuity concerning $\nabla\log q(\sqrt{1-t}\sigma,\mathbf{x})$ with respect to $t$ in Lemma A.4 with $\mathbf{x}$ belonging to a compact set.
A.1 Lipschitz Continuity Regarding Spatial Variable
@@
Let
$p_{\mathbf{x},t,\sigma}(\mathbf{y})=\frac{e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})}{\int_{\mathbb{R}^{d^{*}}}e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}},$
then we have
$\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ $=\frac{\nabla\int_{\mathbb{R}^{d^{*}}}e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}}{\int_{\mathbb{R}^{d^{*}}}e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}}$
$=\frac{\int_{\mathbb{R}^{d^{*}}}\frac{-(\mathbf{x}-\mathbf{y})}{\sigma^{2}(1-t)}e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}}{\int_{\mathbb{R}^{d^{*}}}e^{-\frac{\|{\mathbf{x}-\mathbf{y}}\|^{2}}{2\sigma^{2}(1-t)}}{\widehat{p}}^{*}_{data}(\mathbf{y})\mathrm{d}\mathbf{y}}$
$=-\frac{1}{\sigma^{2}(1-t)}\mathbb{E}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y}),$
and
$\nabla^{2}\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ $=\frac{1}{\sigma^{4}(1-t)^{2}}{\rm{Cov}}_{p_{\mathbf{x},t,\sigma}}(\mathbf{x}-\mathbf{y})-\frac{1}{\sigma^{2}(1-t)}\mathbf{I}_{d}$
$=\frac{1}{\sigma^{4}(1-t)^{2}}{\rm{Cov}}_{p_{\mathbf{x},t,\sigma}}(\mathbf{y})-\frac{1}{\sigma^{2}(1-t)}\mathbf{I}_{d}.$
Under Assumption 4.3, we have
$\mathrm{0}\preccurlyeq\mathrm{Cov}_{p_{\mathbf{x},t,\sigma}}(\mathbf{y})\preccurlyeq d^{*}\mathbf{I}_{d^{*}}.$
Therefore, for any $t\in[0,T]$, we have
$\|\nabla^{2}\log{q}(\sqrt{1-t}\sigma,\mathbf{x})\|\leq{\rm{max}}\left\{\frac{d^{*}}{\sigma^{4}(1-T)^{2}},\frac{1}{\sigma^{2}(1-T)}\right\}.$
This completes the proof.
To prove Lemma 4.3, we firstly introduce two auxiliary lemmas.
Lemma A.1 (Brascamp-Lieb inequality).
Let $\mu(\mathrm{d}\mathbf{x})=\exp(-U(\mathbf{x}))\mathrm{d}\mathbf{x}$ be a probability measure on a convex set $\Omega\subseteq\mathbb{R}^{d}$ whose potential $U:\Omega\rightarrow\mathbb{R}$ is twice continuously differentiable and strictly convex. Then for every locally Lipschitz function $f\in L^{2}(\mu)$,
$\mathrm{Var}_{\mu}(f)\leq\mathbb{E}_{\mu}[\langle\nabla_{\mathbf{x}}f,(\nabla^{2}_{\mathbf{x}}U)^{-1}\nabla_{\mathbf{x}}f].$
Remark A.1.
When applying the functions of the form $f:\mathbf{x}\mapsto\langle\mathbf{x},\mathbf{v}\rangle$ for any $\mathbf{v}\in\mathbb{S}^{d-1}$, the Brascamp-Lieb inequality yields an upper bound of the covariance matrix
$\mathrm{Cov}_{\mu}(\mathbf{x})\preccurlyeq\mathbb{E}_{\mu}[(\nabla_{\mathbf{x}}^{2}U(\mathbf{x}))^{-1}]$
with equality if $\mathbf{x}\sim\mathcal{N}(\mathbf{m},\mathbf{\Sigma})$ with $\mathbf{\Sigma}$ being positive definite.

### v2

@@ 4.2.2 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},{\widehat{p}}^{*}_{data})]$
Let $\mathcal{R}=\lfloor n^{\frac{1}{d^{*}+2\beta}}\rfloor+1\lesssim n^{\frac{1}{d^{*}+2\beta}}$, $T=1-\mathcal{R}^{-C_{T}}$, and $C_{T}=2\beta$, we have
$\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},{\widehat{p}}^{*}_{data})]\lesssim n^{-\frac{\beta}{d^{*}+2\beta}}.$
4.2.3 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},{\widehat{p}}^{*}_{data})]$
By amalgamating the analyses articulated in Theorem 4.6 and Lemma 4.7, we derive an upper bound for $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},{\widehat{p}}^{*}_{data})]$ through the meticulous determination of a suitable score neural network and the optimal stopping time $T$. The end-to-end convergence rate is shown in the following theorem.
@@
Suppose that Assumptions 4.3-4.5 hold, and the score estimator $\widehat{\mathbf{s}}$ defined in (6) is structured as introduced in Theorem 4.5. By choosing $\mathcal{R}$, $m$, $C_{T}$, $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})$ introduced in Theorem 4.5 and Theorem 4.6, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},{\widehat{p}}^{*}_{data})]\lesssim n^{-\frac{\beta}{d^{*}+2\beta}}\log^{\frac{19}{2}}n.$
@@
Theorem 4.8 rigorously establishes the end-to-end theoretical guarantees for the diffusion model within latent space. This is achieved through the seamless integration of score estimation, EM discretization, and early stop techniques. Furthermore, this analysis is independent of the pre-training theory, rendering our theoretical framework applicable to other diffusion models, irrespective of considerations regarding pre-training encoder-decoder structures. This characteristic underscores the independent validity of our theoretical framework. Importantly, the convergence rate $n^{-\frac{\beta}{d^{*}+2\beta}}\log^{\frac{19}{2}}n$ in Theorem 4.8 mitigates the curse of dimensionality of raw data, primarily attributed to the implementation of the encoder component. Additionally, this rate is minimax optimal by ignoring the logarithmic factor.
@@ 4.3 Main Result
Suppose that Assumptions 4.1-4.6 hold, and the score estimator $\widehat{\mathbf{s}}$ defined in (6) is structured as introduced in Theorem 4.5. By choosing $\mathcal{R}$, $m$, $C_{T}$, and $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})$ introduced in Theorem 4.5 and Theorem 4.6, we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T}^{L},p_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{\beta}{d^{*}+2\beta}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\mathcal{M}^{-\frac{1}{2(d+2)}}+\delta_{0}^{1/2}\right),$
where $\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T}^{L}$ denotes the distribution of $\widehat{\boldsymbol{D}}(\widetilde{\mathbf{x}}_{K})$ in Algorithm 1. Moreover, if $\mathcal{M}>n^{\frac{2\beta(d+2)}{d^{*}+2\beta}}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T}^{L},p_{data})]=\widetilde{\mathcal{O}}\left(n^{-\frac{\beta}{d^{*}+2\beta}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\delta_{0}^{1/2}\right).$
@@
In Theorem 4.9, the convergence rates of $\widehat{\boldsymbol{D}}_{\#}\widetilde{\pi}_{T}^{L}$ are established by simultaneously considering pre-training encoder-decoder structures and latent diffusion models. This represents an end-to-end result. Recall that $\mathcal{M}$ represents the sample size of pre-training. This term can be chosen sufficiently large, as the pre-training phase can be manipulated separately by resorting to pre-existing large-scale models. Therefore, we can set $\mathcal{M}>n^{\frac{2\beta(d+2)}{d^{*}+2\beta}}$ to obtain the convergence rate of $\widetilde{\mathcal{O}}\left(n^{-\frac{\beta}{d^{*}+2\beta}}+\epsilon_{p_{data},\widetilde{p}_{data}}+\delta_{0}^{1/2}\right)$. Additionally, if the distribution shift $\epsilon_{p_{data},\widetilde{p}_{data}}$ and pre-training quantity $\delta_{0}^{1/2}$ are both smaller than $n^{-\frac{\beta}{d^{*}+2\beta}}$, then the convergence rate reduces to the optimal order $\widetilde{\mathcal{O}}(n^{-\frac{\beta}{d^{*}+2\beta}})$. This finding circumvents the curse of dimensionality inherent in raw data, thereby demonstrating the effectiveness of latent diffusion models from a theoretical perspective.
@@ 5 Conclusion
In this work, we introduce a novel latent diffusion model rooted in the Schrödinger bridge, facilitated by pre-training an encoder-decoder structure. An SDE, defined over the time interval $[0,1]$, is formulated to effectuate the transformation of the convolution distribution into the encoder target distribution within the latent space. The synergy of pre-training, score estimation, EM discretization, and early stopping enables the derivation of desired samples, manifesting an approximate distribution in accordance with the target distribution. Theoretical analysis encompasses pre-training, score estimation theory, and comprehensive end-to-end convergence rates for the generated samples, quantified in terms of the second-order Wasserstein distance. Our theoretical findings provide solid theoretical guarantees for mainstream diffusion models. The obtained convergence rates predominantly hinge upon the dimension $d^{*}$ of the latent space, achieving optimality and mitigating the curse of dimensionality inherent in raw data. However, the dimension $d^{*}$ is prespecified in practical applications. Therefore, the development of an adaptive approach for selecting it can be regarded as future work. Additionally, extending our theoretical framework to analyze diffusion models based on SDEs, such as OU processes and Langevin SDEs, defined over an infinite time horizon $[0,\infty)$ can also be considered as future work.
@@
In this appendix, we provide the detailed proofs of all lemmas and theorems in this paper.
In Section A, we derive the error bound associated with the pre-training procedure. In Section B, we provide a comprehensive proof of the approximation error. This proof draws inspiration from the approaches outlined in [OAS23, FYWC24]. To complete the proof, in line with the methodology of [OAS23, FYWC24], we present a series of high-probability bounds for $q_{t}(\mathbf{x})$, its derivatives, and $\nabla\log q_{t}(\mathbf{x})$ in Section F.1, and then introduce auxiliary lemmas in Section F.2 to support the approximation of the Hölder class using ReLU networks. In Section C, we establish the statistical error. Furthermore, Section D provides the upper bound for $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[W_{2}(\widetilde{\pi}_{T}^{L},\widehat{p}_{data}^{*})\right]$. Finally, in Section E, we present the proof of our main results.
Appendix A Pre-training
In this section, we prove Lemma 4.1 and Theorem 4.2.
@@ Appendix A Pre-training
We approximate $[\boldsymbol{E}^{*}(\mathbf{y})]_{i}$, $1\leq i\leq d^{*}$, separately. Then, we can obtain an approximation of $\boldsymbol{E}^{*}(\mathbf{y})$ by concatenation. We rescale the input by $\mathbf{y}^{\prime}=\frac{\mathbf{y}+\mathbf{1}}{2}\in[0,1]^{d}$. Such a transformation can be exactly implemented by a single ReLU layer. We define the rescaled function on the transformed input space as $\boldsymbol{E}^{*,\prime}(\mathbf{y}^{\prime}):=\boldsymbol{E}^{*}(2\mathbf{y}^{\prime}-\mathbf{1})$, so that $\boldsymbol{E}^{*,\prime}$ is $2\xi_{\boldsymbol{E}}$-Lipschitz continuous.
We then partition $[0,1]^{d}$ into non-overlapping hypercubes with equal edge length $e_{1}$. $e_{1}$ will be chosen depending on the desired approximation error. We denote $N_{1}=\lceil\frac{1}{e_{1}}\rceil$ and let $\mathbf{v}=[v_{1},v_{2},\cdots,v_{d}]^{\top}\in[N_{1}]^{d}$ be a multi-index. We define $[\overline{\boldsymbol{E}}(\mathbf{y})]_{i}$, $1\leq i\leq d^{*}$, as